OpenStudy (anonymous):

x^4-10x^2+21 divided by x^2-3 =x^2+7?

OpenStudy (anonymous):

oops should be x^2-7

OpenStudy (anonymous):

again, use y=x^2

OpenStudy (turingtest):

you are right tena

OpenStudy (anonymous):

x^4 - 10x^2 + 21 = (x^2-7)(x^2-3). So, the answer is right.